OpenStudy (anonymous):

@Shamim Can you help me with these Physics questions? You were very helpful last time.

OpenStudy (anonymous):

1) Do you know what is the formula for the electric field due to a charged sphere at any point ?

OpenStudy (anonymous):

Yes.. correct. So you know the electric field strength, and you know the distance 'r' so can you find the value of charge now?

OpenStudy (anonymous):

E = electric field strength r = distance q = charge k= constant = 9 X 10^9

OpenStudy (anonymous):

you did something wrong in the math \[q = \frac{Er^2}{k} = \frac{450 \times 0.0625}{9 \times 10^{9}}\]

OpenStudy (anonymous):

there is the 10 power nine in the denominator. How can you get such a big number. ?

OpenStudy (anonymous):

As for 3), you should know that the formula for Power is P=IV, where V is voltage and I is current. Then , you simply plot the values, being P=9*0.39=?W